2. Let X be a continuous random variable with pdf
f(x) = cx^2 , |x| ≤ 1,
0, otherwise,
where the parameter c is constant (with respect to x).
(a) Find the constant c.
(b) Compute the cumulative distribution function F(x) of X.
(c) Use F(x) (from b) to determine P(X ≥ 1/2).
(d) Find E(X) and V (X).
1
Expert's answer
2019-11-11T10:49:53-0500

a) We know that for "f(x)" to be a probability distribution


"\displaystyle\int_{-\infin}^\infin f(x)dx=1"

We integrate "f(x)" with respect to "x," set the result equal to 1 and solve for "c."


"\displaystyle\int_{-\infin}^\infin f(x)dx=\displaystyle\int_{-1}^{1} cx^2dx=c\big[{x^3 \over 3}\big]\begin{array}{c}
 1 \\
 -1
\end{array}="

"=c\big({1^3 \over 3}-{(-1)^3 \over 3}\big)={2 \over 3}c=1"

"c={3 \over 2}"

b) Cumulative distribution function (cdf):


"F(x)=P(X\leq x)=\displaystyle\int_{-\infin}^x f(x)dx"

"\displaystyle\int_{-1}^x {3 \over 2}x^2 dx ={3 \over 2}\big[{x^3 \over 3}\big]\begin{array}{c}
 x \\
 -1
\end{array}={x^3 \over 2}+{1 \over 2}"

"F(x) = \begin{cases}
 0 &\text{if } x<-1 \\
 {x^3 \over 2}+{1 \over 2} &\text{if }-1\leq x<1 \\
 1 &\text{if } x\geq 1
\end{cases}"

c)


"P(X\geq {1 \over 2})=1-P(X< {1 \over 2})=1-P(X\leq {1 \over 2})="

"=1-F({1 \over 2})=1-({({1 \over 2})^3 \over 2}+{1 \over 2})={7 \over 16}"

d)


"\mu=E(X)=\displaystyle\int_{-\infin}^\infin xf(x)dx=\displaystyle\int_{-1}^{1} {3 \over 2}x^3dx="

"={3 \over 2}\big[{x^4 \over 4}\big]\begin{array}{c}
 1 \\
 -1
\end{array}=0"

"Var(X)=\sigma^2=E[(X-\mu)^2]=\displaystyle\int_{-\infin}^\infin (x-0)^2f(x)dx="

"=\displaystyle\int_{-1}^{1} {3 \over 2}x^4dx={3 \over 2}\big[{x^5 \over 5}\big]\begin{array}{c}
 1 \\
 -1
\end{array}={3 \over 5}"


"E(X)=0""Var(X)={3 \over 5}"